OpenStudy (anonymous):

Solve the equation. (If possible) 5/x-5 + 1/x+5 = 2/x^2-25

OpenStudy (anonymous):

Start by multiplying everything by x^2

hero (hero):

Now just wait one minute...I don't think multiplying everything by x^2 is the best way to go.

OpenStudy (anonymous):

ahhh, i was reading the problem differently

hero (hero):

Here's what we should probably do instead: 1. Make sure the denominators are factored: \[\frac{5}{x-5} + \frac{1}{x+5} = \frac{2}{(x+5)(x-5)}\]

OpenStudy (anonymous):

Ok I understand that

OpenStudy (anonymous):

Should I get the denominators to be the same?

hero (hero):

Well, in this case the next step would be to multiply both sides by (x-5)

hero (hero):

What would you end up with if you did that?

OpenStudy (anonymous):

why would i multiply both by x-5?

hero (hero):

Because, if you did that, you would get this: \[5 - \frac{x-5}{x+5} = \frac{2}{x+5}\]

hero (hero):

Now all you would have to do after that is add \(\large \frac{x-5}{x+5}\) to both sides.

OpenStudy (anonymous):

Thank you!

hero (hero):

Oh no...you can't get off that easy. You're not done until you tell me what x equals :P

hero (hero):

If you get it right, I'll give you a medal.